OpenStudy (anonymous):

3x/2 = x + 15 and solve for x

OpenStudy (jackellyn):

Multiply 2 by both sides 3x/2 * 2 = 2(x+15) 3x -2x = 2x +30 -2x x =30

OpenStudy (anonymous):

how did you come to the conclusion of multiplying by 2

OpenStudy (anonymous):